If for a two dimensional potential flow, the velocity potential is given by : φ = 4x(3y - 4), determine the velocity at the point (2, 3). Determine also the value of stream function ψ at the point (2, 3).
1
Expert's answer
2020-12-11T09:56:30-0500

In velocity we can write 


"u=\dfrac{\partial\varphi}{\partial x}, v=\frac{\partial\varphi}{\partial y}"


Given "\varphi=4x(3y-4)"

"u=\dfrac{\partial\varphi}{\partial x}=12y-16, v=\dfrac{\partial\varphi}{\partial y}=12x"

Calculate the velocity at the point (2, 3)


"u=20, v=24"

"|velocity|=\sqrt{u^2+v^2 }=\sqrt{(20)^2+(24)^2}=4\sqrt{61}"

"u=\dfrac{\partial\psi}{\partial y}, v=-\dfrac{\partial\psi}{\partial x}"

"u=\dfrac{\partial\psi}{\partial y}=12y-16"

Integrate with respect to "y"


"\psi=6y^2-16y+g(x)"

"-\frac{\partial\psi}{\partial x}=-\frac{dg}{d x}=12x"

Integrate with respect to "x"


"g(x)=-6x^2+C"

Then


"\psi(x,y)=-6x^2+6y^2-16y+C"